OpenStudy (anonymous):

factor: 1. 81p^2r^2-4 2.-3x^2+75 3. 36r^2+60rp+25p^2

OpenStudy (ajprincess):

u need to factorise all three?

OpenStudy (anonymous):

\[\left\{81 p^2 r^2-4,75-3 x^2,25 p^2+60 p r+36 r^2\right\}\to \]\[\left\{(9 p r-2) (9 p r+2),-3 (x-5) (x+5),(5 p+6 r)^2\right\} \]

OpenStudy (anonymous):

Yes all 3

OpenStudy (ajprincess):

is that clear?

OpenStudy (anonymous):

i'm still a bit confused but I'm trying to figure out. I understand the 1st one. but wouldn't I have to change it since 9 is the a squared 3?

OpenStudy (ajprincess):

vch one u mean? the 1st one?

OpenStudy (anonymous):

I need help one by one

OpenStudy (ajprincess):

1.81p^2r^2-4 This is the of the form a^2-b^2. a^2-b^2=(a-b)(a+b) Here a^2=81p^2r^2 so a=sqrt(81p^2r^2) =9pr b^2=4 so b=2 Is it clear?

OpenStudy (anonymous):

no still confused. Our teacher is teaching us factoring but only wants us to concentrate on factoring squared #'s at the moment. ........ Ok wait I see what you did

OpenStudy (anonymous):

ok so (9pr+2)(9pr-2) .... is that it or do we have to continue factoring because 9 is square of 3?

OpenStudy (ajprincess):

no that's all.

OpenStudy (anonymous):

ok then let me try the others.

OpenStudy (anonymous):

-3x^2+75 common factor is 3 so -3(x^2-25) yes

OpenStudy (anonymous):

36r^2+60rp+25p^2 would be 6r(6r+10rp)+5(5)

OpenStudy (anonymous):

is this correct?

OpenStudy (ajprincess):

-3(x^2-25) is correct bt u can factorise (x^2-25) further.

OpenStudy (anonymous):

how? -3(x-5)(x+5)

OpenStudy (ajprincess):

(x^2-25) (a^2-b^2)=(a-b)(a+b) a^2=x^2 so a=x b^2=25 so b=5 therefore (x^2-25=(x-5)(x+5) Is it clear?

OpenStudy (anonymous):

ok So I got it right? -3(x-5)(x+5)

OpenStudy (anonymous):

are you still there?

OpenStudy (ajprincess):

yes.

OpenStudy (anonymous):

what about this one.... 6r(6r+10rp)+5(5) is this correct?

OpenStudy (ajprincess):

u shud work that ques in this way. 36r^2+60rp+25p^2 =36r^2+30rp+30rp+25p^2 =6r(6r+5p)+5p(6r+5p) =(6r+5p)(6r+5p) =(6r+5p)^2 is it clear?

OpenStudy (anonymous):

Ok let me take a look to make sure I understand

OpenStudy (anonymous):

im confused about this one. Because there is one I had before that looked like this.... 4x^2-12xy+9y^2 = (2x-2xy+3y)(2x+6xy+3y) .... wouldn't I do the same here?

OpenStudy (ajprincess):

. 4x^2-12xy+9y^2 = (2x-3y)^2

OpenStudy (ajprincess):

See when u multiply (2x-2xy+3y)(2x+6xy+3y) if u get 4x^2-12xy+9y^2

OpenStudy (anonymous):

so I did not do that one correct either?

OpenStudy (ajprincess):

yes.

OpenStudy (anonymous):

i understood the first one you broke down. Now Im confused about 4x^2-12xy+9y^2

OpenStudy (ajprincess):

4x^2-12xy+9y^2 =4x^2-6xy-6xy+9y^2 =2x(2x-3y)-3y(2x-3y) =(2x-3y)(2x-3y) =(2x-3y)^2

OpenStudy (anonymous):

This is harder for me to comprehend. It gets frustrating when others in class find it so simple. I understood. Let me try one on my own to check for understanding

OpenStudy (ajprincess):

ha k.

OpenStudy (anonymous):

Im going to attempt 121x^2+22x+1 =121x^2+11x+11x+1 =11x(11x+1)1(11x+1) =(11x+1)(11x+1) =(11x+1)^2

OpenStudy (anonymous):

is this correct?

OpenStudy (ajprincess):

yes. correct.:)

OpenStudy (anonymous):

thank you so very much. Do you have the formula for this...... the way you explained with a & b?

OpenStudy (ajprincess):

u r welcome.:) (a+b)^2=a^2+2ab+b^2